what is the axis of symmetry of this equation
y=2x2+8x-2

Found 2 solutions by drk, stanbon:
Answer by drk(1908)

To find axis of symmetry, we use -b/2a. So,
-b/2a = -8/2(2) = -8/4 = -2
The axis of symmetry is
X = -2.

Answer by stanbon(75887)

what is the axis of symmetry of this equation
y=2x^2+8x-2
2(x^2+4x+? = y+2+?
2(x^2+4x+4) = y+2+8
2(x+2)^2 = y+10
y = 2(x+2)^2 - 10
----------------------
Axis of symmetry: x = -2
==========================
Cheers,
Stan H.